A 1500 kg car traveling at a speed of 10 m/sskids to a halt on wet concrete where μk = 0.60.

How long are the skid marks?

Express your answer using two significant figures.

L=...
Answer #1

initial velocity vi = 10 m/s

final velocity vf = 0


initial kinetic energy ki = (1/2)*m*vi^2

final kinetic energy Kf = (1/2)*m*vf^2

frictional force fk = uk*m*g


work done by frictional force Wf = f*L*cos180 = -uk*m*g*L

from work energy theorem


work = change in kinetic energy

-uk*m*g*L = (1/2)*m*(vf^2-vi^2)

-0.6*9.8*L = (1/2)*(0-10^2)

distance L = 8.5 m <<<<<------ANSWER